OpenStudy (anonymous):

the shorter leg of a 30°- 60°- 90° triangle is 4 how long is the hypotenuse a)2 b)4 √ 2 c)4√ 3 d)8

OpenStudy (anonymous):

@SA5UK3 any idea on how to do this one ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Oh pellet

OpenStudy (anonymous):

Wait no

OpenStudy (anonymous):

I read it wrong

OpenStudy (anonymous):

snap ! of thanks almost marked it wrong

OpenStudy (anonymous):

Lemme do it and I'll let you know. Gimme a few

OpenStudy (anonymous):

okk thanks

OpenStudy (anonymous):

I believe its 8

OpenStudy (anonymous):

Umm no. Damn this is bad. Gimme some more time

OpenStudy (anonymous):

i found the awnser !! lol but thanks !

OpenStudy (anonymous):

its 4√ 2!

OpenStudy (anonymous):

Haha great!

OpenStudy (anonymous):

Now I'm sure the Hyp = 8

OpenStudy (anonymous):

4^2 + (4root3)^2 = 8^2 a^2 + b^2 = c^2

OpenStudy (anonymous):

ohh ok i see oops i guess i did it wrong thanks!

OpenStudy (anonymous):

@SA5UK3 thanks !

OpenStudy (anonymous):

Haha np, sorry I couldn't have gotten to my conclusion faster